# Robust assessment of asymmetric division in colon cancer cells

### v2

@@ II.2 Measuring fluctuations via flow cytometry
Figure 2: Quantification of partitioning noise via population-level measurements. a) Time evolution of CellTrace-VioletTM fluorescence intensity distribution measured in a flow cytometry time course experiment for a population of HCT116 cells. Time progresses from the darker shade of blue to the lightest, spanning from $[0,84]$ hours (bottom line to top). b) Snapshots of the evolution of the distribution of CellTrace-Violet fluorescence intensity measured in a flow cytometry time course experiment for a population of HCT116 cells. Experimental data are represented by the light blue histogram, while the best-fit Gaussian Mixture Model is displayed as lines, with different colors representing different generations. c) Mean (left) and variance (right) of the intensity of the fluorescent markers as a function of generations, normalized to the initial population values. Each replica of the experiment is identified by a different color and point marker’s shape. The experiment showed are conducted on Caco2 cells and the points correspond to the mean values for each generation. Dashed lines represent the best fit according to Equations 10 and 6, respectively. d) Division asymmetry of the $\Pi(f)$ obtained by fitting Equation 10 to data for all experiments and cell lines. The division asymmetry is measured via the percentage of the coefficient of variation.
@@ II.3 Validation of the method via live fluorescence microscopy
The flow cytometry-based protocol combined with the developed theoretical model provides insights into the properties of the cell partition function (see Figure 2). To validate our approach, we performed extensive live time-lapse microscopy experiments to measure the partition fraction of a labelled compound during proliferation, providing a direct measure of the $\Pi(f)$. Specifically, we aimed to follow a growing cell colony over time, detect divisions, evaluate the fluorescence intensity of mother and daughter cells, and calculate the inherited fraction. The detailed experimental procedure is described in the Materials and Methods. Briefly, cells are stained for the cytoplasm and are plated at low density (approximately $10^{3}$ cells/well) on IBIDI cell imaging chambers ($\mu$-Slide 4 and 8 wells). After 24 hours (sufficient time for complete cell adhesion to the slide) the wells are washed, the media is renewed and live imaging acquisition time-lapse is started. Multiple fields are followed in each well, and for each field, we acquire a z-stack bright-field and fluorescence image (a single plane is shown in Figure 3a) at constant intervals (20 min) for 3 days. The fields are manually selected at the beginning of the acquisition based on cell density and fluorescence intensity. The recorded time lapses have been then manually analyzed to identify divisions. A sample outcome of the analysis procedure is shown in Figure 3(b). The total fluorescence of the cells is displayed as a function of time, before and after the division. The mother cell fluorescence decays exponentially over time and at division it halves due to components partitioning between the two daughters. In this specific case, we can observe that the fraction of cytoplasm inherited by the two cells is not equal, indicating an asymmetrical division. To compute the partition fraction $f$, we fit the fluorescence intensity of the daughters’ cells vs time with:
$log(I_{2i}(t))$ $=mt+q_{2i}$ (12)
$log(I_{21+1}(t))$ $=mt+q_{21+1}$ (13)
where $I(t)$ is the fluorescence intensity at time $t$, $m$ accounts for the decaying process and $2i$ and $12i+$ 1 respectively refer to the two daughter cells. We constrain the slope to be the same between the two cells. The fraction of inherited fluorescence, $f$ , is obtained as :
@@
To ensure the highest possible accuracy, we (i) measured the total fluorescence of the mother cell and those of the two daughter cells for at least 2 hours before and after the division event to increase the determination of the fluorescence splitting for each detected mitosis. Moreover, (ii) we removed all dynamics that showed an absolute Pearson correlation value between luminosity and pixel size higher than 0.9 as a function of time (see Supplementary Information). In fact, single cell’s pixel size, which corresponds to the cell size projection on the focal plane, is found to shrink before mitosis and enlarges after division. Fluorescent intensity is instead proportional to the number of stained cytoplasmic proteins, which are expected to remain constant (except for a constant decay of the fluorescence). Therefore, a high correlation between pixel size and fluorescence intensity indicates a high noise-to-signal ratio. Figure 3(c)(bottom) displays the strip plot of the obtained partition fraction, $f$, for Caco2, HCT116, and CCD18Co cells, randomly spread on the y-axis, and their corresponding fitted $\Pi(f)$ distributions (top). Note that via microscopy imaging, it is possible to measure the whole partition distribution and not just its moments; however, hundreds of events are required to get a reliable estimate. Here, we characterize it by fitting the data with a double Gaussian distribution of the form:
@@
that allows for the measure of both $\langle f\rangle$ and $\sigma_{f}^{2}=\sigma^{\prime 2}+f^{2}-f-1/4$.
The fitted distributions clearly show how HCT116 cells are the most symmetric with only one central and narrow peak, while for CCD18Co and CaCO2 cells two asymmetric peaks are visible, associated to higher fluctuations. In Figure 3(d), we compare the outcomes of the flow cytometry vs microscopy measurements. For all three cell lines, the found degrees of asymmetry are statistically consistent between the two approaches.
Figure 3: Quantification of partitioning noise via single-cell measurements. a) Example of a recorded cell colony of HCT116 cells in brightfield (left) and on CTFR fluorescence (right). b) Cell cytoplasm fluorescence intensity as a function of time for a cell before and after division. Dark green circles correspond to the fluorescence intensity of the mother cell up to the division frame, and then to the sum of the daughters’ fluorescence. Lighter green triangles and squares represent the fluorescence intensity of the daughter cells. Solid lines are the linear fit of the points. The intercepts of the linear fit are used to compute the fraction of tagged components inherited by the daughter cells. The time is counted from the start of the experiment. c) (bottom) Strip plot of the distribution of the inherited fraction of cytoplasm for the different cell lines. The points are randomly spread on the y-axis to avoid overlay. (top) Fit of the inherited fraction distribution with the sum of two Gaussians with mean symmetric to $1/2$. d) Comparison of division asymmetry obtained with time-lapse fluorescent microscopy measures (striped bars) with the ones obtained from flow cytometry experiments (plain bars). The flow cytometry bars are obtained as the mean over the multiple conducted experiments.
@@ II.4 Size division bias accounts for cytoplasmic fluctuations
Figure 4: Cytoplasm partition fluctuations vs cells size. a) Behavior of the integral term $\Sigma(N)$ (Eq.19) as a function of the number of dividing elements, $N$, at fixed $\sigma_{N_{i}}=0.8$. Vertical lines mark typical values of cellular elements, like mitochondria. b) Theoretical value of the variance of $\Pi(f)$ in the binomial assumption for different levels of asymmetries and increasing values of $N$. c) Asymmetry of the partitioning distribution $\Pi(f)$ in the binomial limit as measured by the binomial bias, p. d) Sample cases of volume asymmetric division for different cell lines. Images show the overlay of consecutive times during the division dynamic. Time flows from the bottom left to the top right.
In the above sections, we have proposed and validated a method to evaluate fluctuations in the division of adherent cells. The model makes no assumptions about the shape of the partitioning distribution and, in its general form, accounts only for the effect of its variance in shaping the population dynamics. Indeed, we cannot yet distinguish between divisions which are biased, defined by a partitioning distribution with peak values different from $1/2$ (as shown in figure 3(C) for CaCO2 and CCD18Co cells) or symmetric ones (as shown in figure 3(c) for HCT-116 cells) which though have higher fluctuations. Also, apart from measuring it, we want to propose a qualitative interpretation of the obtained level of asymmetry. We start by recalling that our experimental protocol uses dyes that bind a-specifically to cytoplasmic amines. Thus, (i) it is expected to be uniformly distributed in the cellular cytoplasmic space and (ii) the number of labeled cytoplasmic components can be considered large. With this hypothesis, the least complex, model one can assume is a binomial one with parameter $p$, measuring the bias in the process. Via equation  10 we got a direct link between the measurable variance of the population and the second moment of the underlining partition probability distribution, $\Pi(f)$. Henceforth we sought for a relationship between the second moment of $\Pi(f)$ and the parameters of the binomial distribution.
To begin with, we can write the partition distribution for the fraction of inherited component, assuming a level of asymmetry $p$ ($q=1-p$), as:
@@
Recalling that $m_{2i}$ is the number of inherited intra-cellular components of one of the two daughter cells, the partitioning fraction $f$ can be expressed as $f=m_{2i}/N_{i}$ which leads to:
@@
The quantity $\sigma_{N_{i}}$ can be computed from flow cytometry or microscopy measurements. Indeed, if we consider the relation $I=N_{i}F$, where $I$ is the total fluorescence intensity, $N_{i}$ the number of components and $F$ the fluorescence of each component, then $\sigma_{I}^{2}=\sigma_{N_{i}}^{2}$. Since $\sigma_{I}^{2}$ is accessible from the data, so is $\sigma_{N_{i}}$. On the other hand, the values of $\mu$ are not directly accessible, limiting the possibility of an exact analytical computation. However, by computing the value of $\Sigma(N)$ as a function of $N$ for fixed values of $\sigma_{N_{i}}$ (Figure 4(a)), one can see that already for $N=100$ the role of $\Sigma(N)$ is negligible. In general, for the partitioning of components such as cytoplasm which are present in much larger number in the cell, this factor can be neglected. In this regime, the expression for $\sigma_{f}^{2}$ becomes:
@@
In Figure 4(c), we validated the analytical expression to the values obtained via microscopy measurements, showing again good comaptibility of the results. CaCO2 cells confirm to be the most asymmetric among the cell lines and since they are known for their heterogeneity in cell morphology Lea2015 , we explored the hypothesis of shape heterogeneity and cytoplasmic asymmetry to be linked. Figure 4(d) displays three sample cases of division examined through time-lapse fluorescent microscopy following cytoplasm. Different time frames are overlaid to depict the entire dynamic at once, with time progressing from the bottom left to the top right of each image.
@@ III Discussions
Cell division is orchestrated by hundreds of molecular interactions, which are intrinsically stochastic processes bialek2012biophysics . The presence of such stochasticity results in the insurgence of variability among the cells phenotypes sharing the same genome, as it is the case for cancer cells. In this respect, gene expression noise has been extensively characterized Elowitz2002 incorporating cell cycle effects thomas2019intrinsic ; swain2002intrinsic ; chen2004dynamics , as well as the different strategies that cells evolved to regulate the extent of the noise in these channels Osella2011 ; Miotto2019mirna .
@@
Beside such noise sources, proliferating cells are subject to the fluctuations originated by the partition of cellular elements at division. Analysis of this partition noise highlighted that under some circumstances, its contribution exceeds the others in creating heterogeneity Huh2011 ; Soltani2016 . In fact, asymmetrically dividing cells can produce daughters that differ in size, cellular components, and, in turn, fate Chhabra2021 .
@@
In this framework, the origin of such asymmetry may be ascribed to the basal stochasticity of molecular processes taking place at different levels during the division and/or to specific, evolutionary-conserved mechanisms used by cells to control cell fate and generate diversity Sunchu2020 ; evolat . In fact, asymmetric partitioning of components alters the initial counts of molecules in the following cell cycle, which can lead the system to a different phenotypic state moris2016transition ; Miotto2023coll . Here, we show that an accurate determination of partition fluctuations can be obtained via standard flow cytometry measurements of properly marked cell populations. Our approach is easier and faster with respect to the use of microscopy imaging, which requires the acquisition and analysis of extensive recording of the cells, ultimately limiting the reachable statistics. In addition, it is specifically designed to isolate partitioning noise from other sources via the usage of non-endogenous dyes to target cell components.
@@
Although, the need to use live fluorescent dyes to track proliferation surely limits the extent of applicability of our method to markable components, we anticipate that there are numerous contexts in which our approach could give a determining insight. For instance, organelles such as mitochondria, the endoplasmic reticulum, lysosomes, peroxisomes, and centrosomes can be easily marked peruzzi2021asymmetric and their unequal distribution between daughter cells has been observed to provoke functional differences that influence their fate. In this respect, asymmetric division of mitochondria in stem cells is associated with the retention of stemness traits in one daughter cell and differentiation in the other Katajisto2015 and, although the functional consequences are not yet fully determined, to therapeutic resistance in Cancer Stem Cells hitomi2021asymmetric ; garcia2020role . Asymmetric inheritance of lysosomes, together with mitochondria, appears to play a crucial role in determining the fate of human blood stem cells loeffler2022asymmetric . Furthermore our approach could be extended to investigate cell growth dynamics and the origins of cell size homeostasis in adherent cells miotto2023determining ; miotto2024size ; kussell2005phenotypic ; mcgranahan2017clonal ; de2019exploration .
@@
Here, we measured the segregation statistics of the cytoplasm of different colon cell types, finding that they are neatly distinct. CaCo2 cells exhibits the highest degree of cytoplasmic partition fluctuations, while HCT116 division is the one with lowest asymmetry. It is interesting to note that normal colon fibroblasts sit in between, suggesting that the cancer-associated dis-regulation of the normal cell activities leads to the reactivation of asymmetric cell division in a cancer-specific manner. Indeed, CaCo2 cells are known to display a markedly heterogeneous distribution of phenotypes Lea2015 . Notably, comparing the degree of asymmetry in the division of the cytoplasm in the three different analysed cell types, we found that microscopy-derived fluctuations are systematically higher than those measured by flow cytometry. While this could be linked to the different used protocols, it seems more probable that the difference is due to a systematic overestimation of the fluctuations measured by microscopy due to its lower statistics.
@@
Finally, to get insights on the possible mechanisms behind the observed asymmetries, we looked for specific shape of the partition distribution. Following a maximum entropy approach bialek2012biophysics ; Miotto2018 ; TOLOMEO , we started by considering the distribution that employs the fewest assumptions about the system, i.e., we consider independent segregation, which directly maps into binomial partitioning. Indeed, this is a common assumption in segregation models that traces back to the pioneering works of Berk berg1978model and Rigney rigney1979stochastic . Notably, our results indicated biased segregations with colon cancer cells displaying different levels of biases. To explain their origin, we looked at relative sizes between the daughter cells right after division. The found correlation between daughter sizes and the fractions of inherited fluorescence vouches size as the origin of the observed biased segregation.
In conclusion, by combining experimental data with statistical modeling, we show how it is possible to use flow cytometry data to extract reliable estimates of the strength of fluctuations during cell division. Our approach has the potential to be applied to different cell types where a quantification of the level of division asymmetries daughter cells can experience may provide new insights into the mechanisms of asymmetric cell division and its role in cancer heterogeneity and plasticity.

### v3

@@ II.2 Measuring fluctuations via flow cytometry
Figure 2: Quantification of partitioning noise via population-level measurements. a) Time evolution of CellTrace-VioletTM fluorescence intensity distribution measured in a flow cytometry time course experiment for a population of HCT116 cells. Time progresses from the darkest shade of blue to the lightest, spanning from $[0,84]$ hours (bottom line to top). b) Snapshots of the evolution of the distribution of CellTrace-Violet fluorescence intensity measured in a flow cytometry time course experiment for a population of HCT116 cells. Experimental data are represented by the light blue histogram, while the best-fit Gaussian Mixture Model is displayed as lines, with different colors representing different generations. c) Mean (left) and variance (right) of the intensity of the fluorescent markers as a function of generations, normalized to the initial population values. Each replica of the experiment is identified by a different color and a point marker’s shape. The experiments are conducted on Caco2 cells, and the points correspond to the mean values for each generation. Dashed lines represent the best fit according to Equations 10 and 6, respectively. d) Division asymmetry of the $\Pi(f)$ obtained by fitting Equation 10 to data for all experiments and cell lines. The division asymmetry is measured via the percentage of the coefficient of variation.
@@ II.3 Validation of the method via live fluorescence microscopy
The flow cytometry-based protocol combined with the developed theoretical model provides insights into the properties of the cell partition function (see Figure 2). To validate our approach, we performed extensive live time-lapse microscopy experiments to measure the partition fraction of a labelled compound during proliferation, providing a direct measure of the $\Pi(f)$. Specifically, we aimed to follow a growing cell colony over time, detect divisions, evaluate the fluorescence intensity of mother and daughter cells, and calculate the inherited fraction. The detailed experimental procedure is described in the Materials and Methods. Briefly, cells are stained for the cytoplasm and are plated at low density (approximately $10^{3}$ cells/well) on IBIDI cell imaging chambers ($\mu$-Slide 4 and 8 wells). After 24 hours (sufficient time for complete cell adhesion to the slide), the wells are washed, the media is renewed, and live imaging acquisition time-lapse is started. Multiple fields are followed in each well, and for each field, we acquire a z-stack bright-field and fluorescence image (a single plane is shown in Figure 3a) at constant intervals (20 min) for 3 days. The fields are manually selected at the beginning of the acquisition based on cell density and fluorescence intensity. The recorded time lapses have been manually analyzed to identify divisions. A sample outcome of the analysis procedure is shown in Figure 3(b). The total fluorescence of the cells is displayed as a function of time, before and after the division. The mother cell fluorescence decays exponentially over time, and at division, it halves due to components partitioning between the two daughters. In this specific case, we can observe that the fraction of cytoplasm inherited by the two cells is not equal, indicating an asymmetrical division. To compute the partition fraction $f$, we fit the fluorescence intensity of the daughters’ cells vs time with $log(I_{2i}(t))=mt+q_{2i}$ and $log(I_{21+1}(t))=mt+q_{21+1}$, where $I(t)$ is the fluorescence intensity at time $t$, $m$ accounts for the decaying process and $2i$ and $2i+1$ respectively refer to the two daughter cells. We constrain the slope to be the same between the two cells. The fraction of inherited fluorescence, $f$ , is obtained as :
@@
To ensure the highest possible accuracy, we (i) measured the total fluorescence of the mother cell and those of the two daughter cells for at least 2 hours before and after the division event to increase the determination of the fluorescence splitting for each detected mitosis. Moreover, (ii) we removed all dynamics that showed an absolute Pearson correlation value between luminosity and pixel size higher than 0.9 as a function of time (see Supplementary Information). In fact, single cell’s pixel size, which corresponds to the cell size projection on the focal plane, is found to shrink before mitosis and to enlarge after division. Fluorescent intensity is instead proportional to the number of stained cytoplasmic proteins, which are expected to remain constant (except for a constant decay of the fluorescence). Therefore, a high correlation between pixel size and fluorescence intensity indicates a high noise-to-signal ratio. Figure 3(c)(bottom) displays the strip plot of the obtained partition fraction, $f$, for Caco2, HCT116, and CCD18Co cells, randomly spread on the y-axis, and their corresponding fitted $\Pi(f)$ distributions (top). Note that via microscopy imaging, it is possible to measure the whole partition distribution and not just its moments; however, hundreds of events are required to get a reliable estimate. Here, we characterize it by fitting the data with a double Gaussian distribution of the form:
@@
that allows for the measure of both $\langle f\rangle$ and $\sigma_{f}^{2}=\sigma^{\prime 2}+f^{2}-f+1/4$.
The fitted distributions clearly show how HCT116 cells are the most symmetric, with only one central and narrow peak, while for CCD18Co and CaCO2 cells, two asymmetric peaks are visible, associated with higher fluctuations. In Figure 3(d), we compare the outcomes of the flow cytometry vs microscopy measurements. For all three cell lines, the found degrees of asymmetry are statistically consistent between the two approaches.
Figure 3: Quantification of partitioning noise via single-cell measurements. a) Example of a recorded cell colony of HCT116 cells in brightfield (left) and on CTFR fluorescence (right). b) Cell cytoplasm fluorescence intensity as a function of time for a cell before and after division. Dark green circles correspond to the fluorescence intensity of the mother cell up to the division frame, and then to the sum of the daughters’ fluorescence. Lighter green triangles and squares represent the fluorescence intensity of the daughter cells. Solid lines are the linear fit of the points. The intercepts of the linear fit are used to compute the fraction of tagged components inherited by the daughter cells. The time is counted from the start of the experiment. c) (bottom) Strip plot of the distribution of the inherited fraction of cytoplasm for the different cell lines. The points are randomly spread on the y-axis to avoid overlay. (top) Fit of the inherited fraction distribution with the sum of two Gaussians with a mean symmetric to $1/2$. d) Comparison of division asymmetry obtained with time-lapse fluorescent microscopy measures (striped bars) with the ones obtained from flow cytometry experiments (plain bars). The flow cytometry bars are obtained as the mean over the multiple conducted experiments.
@@ II.4 Size division bias accounts for cytoplasmic fluctuations
Figure 4: Cytoplasm partition fluctuations vs cells size. a) Behavior of the integral term $\Sigma(N)$ (Eq.17) as a function of the number of dividing elements, NN, at fixed $\sigma_{N_{i}}=0.8$. Vertical lines mark typical values of cellular elements, like mitochondria. b) Theoretical value of the variance of $\Pi(f)$ in the binomial assumption for different levels of asymmetries and increasing values of $N$. c) Asymmetry of the partitioning distribution $\Pi(f)$ in the binomial limit as measured by the binomial bias, p. d) Sample cases of volume asymmetric division for different cell lines. Images show the overlay of consecutive times during the division dynamic. Time flows from the bottom left to the top right.
In the above sections, we have proposed and validated a method to evaluate fluctuations in the division of adherent cells. The model makes no assumptions about the shape of the partitioning distribution and, in its general form, accounts only for the effect of its variance in shaping the population dynamics. Indeed, we cannot yet distinguish between divisions which are biased, defined by a partitioning distribution with peak values different from $1/2$ (as shown in Figure 3(C) for CaCO2 and CCD18Co cells) or symmetric ones (as shown in Figure 3(c) for HCT-116 cells) which though have higher fluctuations. Also, apart from measuring it, we want to propose a qualitative interpretation of the obtained level of asymmetry. We start by recalling that our experimental protocol uses dyes that bind a-specifically to cytoplasmic amines. Thus, (i) it is expected to be uniformly distributed in the cellular cytoplasmic space and (ii) the number of labeled cytoplasmic components can be considered large. With this hypothesis, the least complex model one can assume is a binomial one with parameter $p$, measuring the bias in the process. Via equation  10, we got a direct link between the measurable variance of the population and the second moment of the underlying partition probability distribution, $\Pi(f)$. Henceforth, we sought for a relationship between the second moment of $\Pi(f)$ and the parameters of the binomial distribution.
To begin with, we can write the partition distribution for the fraction of inherited component, assuming a level of asymmetry $p$ (q=1pq=1-p), as:
@@
Recalling that $m_{2i}$ is the number of inherited intracellular components of one of the two daughter cells, the partitioning fraction $f$ can be expressed as $f=m_{2i}/N_{i}$, which leads to:
@@
The quantity $\sigma_{N_{i}}$ can be computed from flow cytometry or microscopy measurements. Indeed, if we consider the relation $I=N_{i}F$, where $I$ is the total fluorescence intensity, $N_{i}$ the number of components, and $F$ the fluorescence of each component, then $\sigma_{I}^{2}=\sigma_{N_{i}}^{2}$. Since $\sigma_{I}^{2}$ is accessible from the data, so is $\sigma_{N_{i}}$. On the other hand, the values of $\mu$ are not directly accessible, limiting the possibility of an exact analytical computation. However, by computing the value of $\Sigma(N)$ as a function of $N$ for fixed values of $\sigma_{N_{i}}$ (Figure 4(a)), one can see that already for $N=100$ the role of $\Sigma(N)$ is negligible. In general, for the partitioning of components such as cytoplasm, which are present in much larger numbers in the cell, this factor can be neglected. In this regime, the expression for $\sigma_{f}^{2}$ becomes:
@@
In Figure 4(c), we validated the analytical expression against the values obtained via microscopy measurements, showing again good compatibility of the results. CaCO2 cells confirm to be the most asymmetric among the cell lines, and since they are known for their heterogeneity in cell morphology Lea2015 , we explored the hypothesis that shape heterogeneity and cytoplasmic asymmetry are linked. Figure 4(d) displays three sample cases of division examined through time-lapse fluorescent microscopy following cytoplasm. Different time frames are overlaid to depict the entire dynamic at once, with time progressing from the bottom left to the top right of each image.
@@ III Discussions
Cell division is orchestrated by hundreds of molecular interactions, which are intrinsically stochastic processes bialek2012biophysics . The presence of such stochasticity results in the insurgence of variability among the cell phenotypes sharing the same genome, as is the case for cancer cells. In this respect, gene expression noise has been extensively characterized Elowitz2002 , incorporating cell cycle effects thomas2019intrinsic ; swain2002intrinsic ; chen2004dynamics , as well as the different strategies that cells evolved to regulate the extent of the noise in these channels Osella2011 ; Miotto2019mirna .
@@
Besides such noise sources, proliferating cells are subject to the fluctuations originating from the partition of cellular elements at division. Analysis of this partition noise highlighted that under some circumstances, its contribution exceeds the others in creating heterogeneity Huh2011 ; Soltani2016 . In fact, asymmetrically dividing cells can produce daughters that differ in size, cellular components, and, in turn, fate Chhabra2021 .
@@
In this framework, the origin of such asymmetry may be ascribed to the basal stochasticity of molecular processes taking place at different levels during the division and/or to specific, evolutionarily conserved mechanisms used by cells to control cell fate and generate diversity Sunchu2020 ; evolat . In fact, asymmetric partitioning of components alters the initial counts of molecules in the following cell cycle, which can lead the system to a different phenotypic state moris2016transition ; Miotto2023coll . Here, we show that an accurate determination of partition fluctuations can be obtained via standard flow cytometry measurements of properly marked cell populations. Our approach is easier and faster with respect to the use of microscopy imaging, which requires the acquisition and analysis of extensive recordings of the cells, ultimately limiting the reachable statistics. In addition, it is specifically designed to isolate partitioning noise from other sources via the usage of non-endogenous dyes to target cell components.
@@
Although the need to use live fluorescent dyes to track proliferation surely limits the extent of applicability of our method to markable components, we anticipate that there are numerous contexts in which our approach could give a determining insight. For instance, organelles such as mitochondria, the endoplasmic reticulum, lysosomes, peroxisomes, and centrosomes can be easily marked peruzzi2021asymmetric ; Miotto2025phyl and their unequal distribution between daughter cells has been observed to provoke functional differences that influence their fate. In this respect, asymmetric division of mitochondria in stem cells is associated with the retention of stemness traits in one daughter cell and differentiation in the other Katajisto2015 and, although the functional consequences are not yet fully determined, with therapeutic resistance in Cancer Stem Cells hitomi2021asymmetric ; garcia2020role . Asymmetric inheritance of lysosomes, together with mitochondria, appears to play a crucial role in determining the fate of human blood stem cells loeffler2022asymmetric . Furthermore, our approach could be extended to investigate cell growth dynamics and the origins of cell size homeostasis in adherent cells miotto2023determining ; miotto2024size ; kussell2005phenotypic ; mcgranahan2017clonal ; de2019exploration .
@@
Here, we measured the segregation statistics of the cytoplasm of different colon cell types, finding that they are neatly distinct. CaCo2 cells exhibit the highest degree of cytoplasmic partition fluctuations, while HCT116 division is the one with the lowest asymmetry. It is interesting to note that normal colon fibroblasts sit in between, suggesting that the cancer-associated deregulation of the normal cell activities leads to the reactivation of asymmetric cell division in a cancer-specific manner. Indeed, CaCo2 cells are known to display a markedly heterogeneous distribution of phenotypes Lea2015 . Notably, comparing the degree of asymmetry in the division of the cytoplasm in the three different analysed cell types, we found that microscopy-derived fluctuations are systematically higher than those measured by flow cytometry. While this could be linked to the different protocols used, it seems more probable that the difference is due to a systematic overestimation of the fluctuations measured by microscopy. Indeed, we identified at least two possible origins of such bias: fluctuations in the fluorescence intensity measures and errors in the segmentation protocol. Regarding the first one, we demonstrated its effect through simple simulations that mimicked the microscopy analysis. While for the second, relying on manual segmentation not only reduces the data but might also introduce biased cell identification errors. Both these effects become important when data filtering is only partially allowed due to the low available statistics. A discussion on this topic is proposed in the Supplementary Information.
@@
Finally, to get insights into the possible mechanisms behind the observed asymmetries, we looked for a specific shape of the partition distribution. Following a maximum entropy approach bialek2012biophysics ; Miotto2018 ; TOLOMEO , we started by considering the distribution that employs the fewest assumptions about the system, i.e., we consider independent segregation, which directly maps into binomial partitioning. Indeed, this is a common assumption in segregation models that traces back to the pioneering works of Berk berg1978model and Rigney rigney1979stochastic . Notably, our results indicated biased segregations with colon cancer cells displaying different levels of biases. To explain their origin, we looked at the relative sizes of the daughter cells right after division. The found correlation between daughter sizes and the fractions of inherited fluorescence vouches size as the origin of the observed biased segregation.
In conclusion, by combining experimental data with statistical modeling, we show how it is possible to use flow cytometry data to extract reliable estimates of the strength of fluctuations during cell division. Our approach has the potential to be applied to different cell types, where a quantification of the level of division asymmetries that daughter cells can experience may provide new insights into the mechanisms of asymmetric cell division and its role in cancer heterogeneity and plasticity.